Sanjay is nine times the age of his son, Anand. In five years time he will be four time the age of Anand. How old are they now
Found 2 solutions by ankor@dixie-net.com, Khom :
Answer by ankor@dixie-net.com(22740)

Let s = S's present age
let a = A's age
:
Write an equation for each statement
:
Sanjay is nine times the age of his son, Anand.
s = 9a
In five years time he will be four time the age of Anand.
s + 5 = 4(a + 5)
s + 5 = 4a + 20
s = 4a + 20 - 5
s = 4a + 15
replace s with 9a (from the 1st statement)
9a = 4a + 15
9a - 4a = 15
5a = 15
a = 15/5
a = 3 yrs is A's age
then
9(3) = 27 yrs is S's age
:
:
See if this checks out in the statement
" In five years time he will be four time the age of Anand."
27 + 5 = 4(3+5)
32 = 4(8)


Answer by Khom (1)

Let present age of Sanjay be 'x'
Present age of his son Anand be 'y'
So, by question
X=9y..... (i)
After 5 years, x+5=4(y+5)..... (ii)
Solving eq^ns (i) and (ii)
We get
.
.
.
.
X =27 years i.e. present age of sanjay
y=3 years i.e. present age of anand